OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = x^2 + 8x + 12.

OpenStudy (anonymous):

the x coordinate of the vertex can be obtained by x = -b/(2a) , where a, b are the coefficients in the standard way of writing the quadratic equation...

OpenStudy (anonymous):

x=-b/2a a=1 b=8 c=12

OpenStudy (anonymous):

ax^2 + bx + c = 0...

OpenStudy (anonymous):

For any parabola of the form \(ax^2+bx+c=0 \) the vertex will be at \( \large \left(-\frac b {2a} , -\frac {b^2-4ac}{4a}\right) \).

OpenStudy (anonymous):

so -(8)/2*1 = -4

OpenStudy (anonymous):

now plug that back in to the quadratic get the value of y if you need it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(-4, 12) (-4, -4) (0, -6)

OpenStudy (anonymous):

now let me c

OpenStudy (anonymous):

have you got it :D

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

\[y=x^2+8x+12 = -4\] wherever you see x put -4 and do the sum

OpenStudy (anonymous):

y = (-4)^2 + 8x(-4) + 12 = ??

OpenStudy (anonymous):

y = (-4)^2 + 8(-4) + 12 = ??

OpenStudy (anonymous):

-36

OpenStudy (anonymous):

so -4x-4 =16 -4x8 = -32 then add 12 16-32+12 =-4

OpenStudy (anonymous):

so the vertex is at (-4,-4)

OpenStudy (anonymous):

ok thank you